OpenStudy (anonymous):

Help solve (3-a)(b+c) 9/x-5/x 52;x/8=6

OpenStudy (anonymous):

are there three questions or ... ? \[\LARGE (3-a)(b+c)=3b+3c-ab-ac\] \[\huge \frac{a}{x}\pm \frac bx=\frac{a+b}{x}\] and you have.. \[\LARGE \frac9x-\frac5x=?\] can you do it ??

OpenStudy (anonymous):

yeh the third one is 52;x/8=6

OpenStudy (anonymous):

... sorry I made a mistake it's..\[\huge \frac{a}{x}\pm\frac{b}{x}=\frac{a\pm b}{x}\]

OpenStudy (anonymous):

Thanks so much.....would that be 4/x???

OpenStudy (anonymous):

\[\LARGE \frac{52x}{8}=6\] .... use cross multiply and you'll get \[\LARGE 52x=6\cdot 8\] \[\LARGE x=\frac{6\cdot 8}{52}\] I'll leave you to finish it...

OpenStudy (anonymous):

yes option b is 4/x ... well done ;)

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

Glad to help ;) ... You're welcome :)

OpenStudy (anonymous):

x=48/52 would that be right

OpenStudy (anonymous):

now simplify it ... 48:2 and 52:2 you'll get: \[\LARGE x=\frac{48}{52}\Longrightarrow x=\frac{24}{26}\] simplify by 2 again... (or other numbers if it's necessary ... till there's no more to simplify ;) ) can you do it ?

OpenStudy (anonymous):

x=12/13

OpenStudy (anonymous):

that's correct ... well done ! ;)

OpenStudy (anonymous):

Thank you for all your help. I struggle bad in algebra, never had to take it til now.

OpenStudy (anonymous):

Glad to help... and again, you're welcome !